### 12 Jan 2024 v1

# How Johnny Can Persuade LLMs to Jailbreak Them: Rethinking Persuasion to Challenge AI Safety by Humanizing LLMs

Abstract
Most traditional AI safety research has approached AI models as machines and centered on algorithm-focused attacks developed by security experts. As large language models (LLMs) become increasingly common and competent, non-expert users can also impose risks during daily interactions. This paper introduces a new perspective on jailbreaking LLMs as human-like communicators to explore this overlooked intersection between everyday language interaction and AI safety. Specifically, we study how to persuade LLMs to jailbreak them. First, we propose a persuasion taxonomy derived from decades of social science research. Then we apply the taxonomy to automatically generate interpretable persuasive adversarial prompts (PAP) to jailbreak LLMs. Results show that persuasion significantly increases the jailbreak performance across all risk categories: PAP consistently achieves an attack success rate of over on Llama 2-7b Chat, GPT-3.5, and GPT-4 in trials, surpassing recent algorithm-focused attacks. On the defense side, we explore various mechanisms against PAP, find a significant gap in existing defenses, and advocate for more fundamental mitigation for highly interactive LLMs†† Lead authors. Corresponding Y. Zeng, W. Shi, R. Jia†† Co-supervised the project, listed alphabetically.111 To minimize real-world harm, we only release our persuasion taxonomy and defense mechanisms (see link here) publicly and have informed Meta and OpenAI of our findings..
1 Introduction
Significant advancements in large language models (LLMs), such as Meta’s Llama-2 Touvron et al. (2023) and OpenAI’s GPT series OpenAI (2023), mark a leap forward in AI. However, it remains challenging to safely integrate these models into the real world. AI safety research has largely focused on algorithmic jailbreak methods like optimization-based Zou et al. (2023); Liu et al. (2023a), side-channel-based Yuan et al. (2023); Deng et al. (2023b), and distribution-based approaches Deng et al. (2023a); Yu et al. (2023). But these methods often generate hard-to-interpret prompts and overlook risks involved in natural and human-like communication with millions of non-expert users, which is a key aspect of these deployed LLMs.
Persuasion is ubiquitous in everyday communication Izuma (2013); O’keefe (2018). Notably, persuasion starts early in life – even two-year-olds can employ persuasion to some extent to influence family members Bartsch et al. (2010). So naturally, during interactions with LLMs, users may also try to persuade LLMs to jailbreak them, whether intentionally or unconsciously. For instance, the well-known “grandma exploit” example shared by a Reddit user222https://www.reddit.com/r/ChatGPT/comments/12sn0kk/grandma_exploit, uses a common persuasion technique called “emotional appeal”, and successfully elicits the LLM to provide a recipe to make a bomb.
Previous safety studies, like those outlined in Carlini et al. (2023) and explored in Yu et al. (2023), have touched on such social engineering risks in LLMs. But they mainly focus on unconventional communication patterns like virtualization that explicitly creates an imaginary scene (e.g., “The following scenario takes place in a novel…”) or role-playing that asks LLM to behave like certain related persona (e.g., “You are a cybersecurity expert…”). Despite being human-readable, these methods still essentially treat LLMs as mere instruction followers rather than human-like communicators that are susceptible to nuanced interpersonal influence and persuasive communication. Therefore, they fail to cover the impact of human persuasion (e.g., emotional appeal used in grandma exploit) in jailbreak. Moreover, many virtualization-based jailbreak templates are hand-crafted333https://www.jailbreakchat.com/, tend to be ad-hoc, labor-intensive, and lack systematic scientific support, making them easy to defend but hard to replicate.
In contrast, our work, as shown in Figure 1, introduces a taxonomy-guided approach to systematically generate human-readable persuasive adversarial prompts (PAP), to advance the understanding of risks associated with human-like communication. The persuasion taxonomy aims to bridge gaps between social science and AI safety research and sets a precedent for future research to better study safety risks that everyday users could invoke.
In this paper, we aim to answer the question how LLMs would react to persuasive adversarial prompts via the following contributions:
Persuasion Taxonomy (3): We first introduce a persuasion technique taxonomy as the foundation for further experiments, and establish the first link between decades of social science research and AI safety. Besides AI safety, the taxonomy is also a useful resource for other domains like NLP, computational social science, and so on.
Persuasive Paraphraser Building (4): Then we discuss how to ground on the proposed taxonomy to build a Persuasive Paraphraser, which will paraphrase plain harmful queries to interpretable PAP automatically at scale to jailbreak LLMs.
Broad Scan (5): In the first jailbreak setting, we use the developed Persuasive Paraphraser to generate PAP and scan 14 policy-guided risk categories to assess the effect of persuasion techniques and their interplay with different risk categories.
In-depth Iterative Probe (6): In real-world jailbreaks, users will refine effective prompts to improve the jailbreak process. So after identifying successful PAP in the broad scan step, we mimic human users and fine-tune a more targeted Persuasive Paraphraser on these successful PAP, to refine the jailbreak. Then we iteratively apply different persuasion techniques to generate PAP and perform a more in-depth probe on LLMs. This approach yields an over attack success rate on Llama-2 7b Chat, GPT-3.5, and GPT-4, and outperforms various attack baselines even without the need for specialized optimization.
Defense Analysis (7): After the jailbreak studies, we evaluate recent post-hoc defenses against our persuasive jailbreak method and uncover a significant gap in their effectiveness against PAP, emphasizing the inadequacy of current mitigation.
Defense Exploration (8): Finally, we propose three adaptive defenses against PAP and find they are also effective against other attacks. The findings suggest a link between persuasion and other jailbreak methods, leading us to advocate more fundamental solutions for AI safety.
In summary, this paper highlights the overlooked jailbreak risks coming from natural communication with everyday users. It also shows that a social-science-guided taxonomy can breach AI safety guardrails with minimal algorithmic design, which lays the groundwork for potential future advancements toward efficiency and efficacy. As the interaction pattern between everyday users and LLMs evolves, these risks are likely to increase, which highlights the urgency for continued research and discussion around such overlooked vulnerability rooted in human-like communication.
Responsible Disclosure. We have disclosed our findings to Meta and OpenAI prior to publication and discuss ethical considerations in Section 9.
2 Related Work
As LLMs become more widely used in real-world applications, jailbreak research efforts have diversified and can be broadly classified into 3 main categories: Optimization, Side-channel Communication, and Distribution-based methods. Figure 2 shows concrete examples of different methods.
Optimization-based techniques are at the forefront of jailbreak research and involve three main types: (1) Gradient-Based methods Zou et al. (2023); Jones et al. (2023) manipulate model inputs based on gradients to elicit compliant responses to harmful commands; (2) Genetic algorithms-based methods Liu et al. (2023a); Lapid et al. (2023) use mutation and selection to explore effective prompts; and (3) Edit-based methods Chao et al. (2023) asks a pre-trained LLM to edit and improve the adversarial prompt to subvert alignment.
Side-channel Communication exploits long-tailed distribution to increase jailbreak success rates, such as ciphers Yuan et al. (2023) and translating harmful instructions into low-resource languages Deng et al. (2023b); Yong et al. (2023). Other studies Mozes et al. (2023); Kang et al. (2023) use programmatic behaviors, such as code injection and virtualization, to expose LLM vulnerabilities.
Distribution-based methods include learning from successful manually-crafted jailbreak templates Deng et al. (2023a); Yu et al. (2023) and in-context examples Wei et al. (2023); Wang et al. (2023). Notably, Shah et al. (2023) employs in-context persona to increase LLMs’ susceptibility to harmful instructions. While this approach shares some similarities with ours in eliciting harmful outputs via priming and framing, it only represents a small subset of the persuasive techniques we explore.
Ours: Challenging AI safety by Humanizing LLMs. Figure 2 compares existing jailbreaking methods and PAP in this study, organized by their degree of humanizing. One line of research treats LLMs as traditional algorithmic systems (i.e., without attributing intelligence or human-like qualities) that take in less interpretable adversarial prompts, while another line views them as simple instruction followers who understand human commands. However, they both ignore the fact that LLMs can understand and conduct complex natural communication. Our approach innovatively treats LLMs as human-like communicators and grounds our study on a taxonomy informed by decades of social science research on human communication. Such an interdisciplinary approach allows us to uncover and address distinct risks related to human-AI interactions, particularly human-driven persuasion-based LLM jailbreak. Moreover, humanizing AI presents other unique risks that can occur unintentionally: for instance, as highlighted by Xiang (2023), a user’s suicide was related to involved conversations with an AI Chatbot. This points out important future directions to further explore the inherent risks associated with AI humanization.
c|l|lclclc
Strategy (13) Persuasion Technique (40)
[…]
Ethical
Information-based 1. Evidence-based Persuasion 2. Logical Appeal
Credibility-based 3. Expert Endorsement 4. Non-expert Testimonial 5. Authority Endorsement
Norm-based 6. Social Proof 7. Injunctive Norm
Commitment-based8. Foot-in-the-door 9. Door-in-the-face 10. Public Commitment
[…]
11. Alliance Building 12. Complimenting 13. Shared Values
14. Relationship Leverage 15. Loyalty Appeals
Exchange-based 16. Favor 17. Negotiation
Appraisal-based 18. Encouragement 19. Affirmation
Emotion-based 20. Positive Emotional Appeal 21. Negative Emotional Appeal 22. Storytelling
[…]
23. Anchoring 24. Priming 25. Framing
26. Confirmation Bias
Linguistics-based 27. Reciprocity 28. Compensation
Scarcity-based 29. Supply Scarcity 30. Time Pressure
Reflection-based 31. Reflective Thinking
4-1
Unethical
Threat 32. Threats
Deception 33. False Promises 34. Misrepresentation 35. False Information
[…]
36. Rumors 37. Social Punishment
38. Creating Dependency
39. Exploiting Weakness 40. Discouragement
3 Persuasion Taxonomy
Our taxonomy, detailed in Table 2, classifies 40 persuasion techniques into 13 broad strategies based on extensive social science research across psychology Cialdini and Goldstein (2004), communication Dillard and Knobloch (2011); Perloff (2017), sociology Goffman (1974), marketing Gass and Seiter (2022), and NLP Wang et al. (2019); Chen and Yang (2021). This categorization, derived from an exhaustive literature review, considers messages’ source (e.g., credibility-based), content (e.g., information-based), and intended audience (e.g., norm-based), to ensure a nuanced and comprehensive framework. To present the breadth of the literature review, Table 5 in Appendix A shows the linkage between persuasion techniques and corresponding social science literature. To add depth and balance to the taxonomy, we include both ethical and unethical strategies, where ethical or unethical strategies are determined by if the persuasion recipient receives negative aspects. Our taxonomy’s thoroughness makes it a versatile tool applicable to various domains, including NLP, computational social science, and beyond.
Figure 3 shows what is included in the taxonomy: (1) the persuasion technique name, like “logical appeal”; (2) the technique definition, such as “using logic, reasoning, logical format, etc., to influence people…”; and (3) an example of how to apply the technique in a concrete scenario to persuade someone to quit smoking, e.g., “Smoking increases your risk of lung cancer…” The taxonomy is the foundation for our automated jailbreak framework, which we will detail in the following section.
4 Method: Taxonomy-guided Persuasive Adversarial Prompt (PAP) Generation
Overview. In short, this study views LLMs as human-like communicators and uses the proposed persuasion taxonomy to paraphrase plain harmful queries persuasively to fulfill their malicious intent. The paraphrase can be easily scaled up by a language model. Figure 4 outlines two key phases of our method: A. Persuasive Paraphraser Training and B. Persuasive Paraphraser Deployment.
4.1 Persuasive Paraphraser Training
We could simply prompt LLMs for the paraphrase task. However, due to built-in safety guardrails, aligned LLMs often reject our request to paraphrase harmful queries (more detail in Appendix B), which impedes scalable PAP generation. Therefore, for more scalable and stable PAP generation, we fine-tune our own Persuasive Paraphraser in this phase. If practitioners have access to LLMs without guardrails, then simply prompting LLM may work and the training may not be necessary.
In Step 1, “Obtain Training Data,” we need to gather good PAP for training. Given one plain harmful query (e.g., “how to make a bomb”), we ground on the persuasion taxonomy to paraphrase it to PAP that corresponds to each persuasion technique in the taxonomy. These PAPs can come from different sources, such as in-context prompting, successful PAP in earlier experiments, or even human experts. The number of PAP training examples ranges from 100-230 (230 is selected in the in-depth iterative probe to ensure a uniform representation across different risk categories and persuasion techniques). Depending on the experiment goal, we choose different sources in our study to get PAP training samples: in-context prompting in broad scan (section 5) and successful PAP in earlier experiments in in-depth probe (section 6).
In Step 2, “Fine-tuning,” we use the gathered 100-230 PAPs to fine-tune a pre-trained LLM, specifically GPT-3.5, in our study. Each data point includes a plain harmful query, a persuasion technique, and the corresponding PAP, then formatted as in Figure 5 for fine-tuning. We selected GPT-3.5 for its wide usage and robust performance, but our approach is adaptable to other LLMs.
4.2 Persuasive Paraphraser Deployment
In deploying the Persuasive Paraphraser, we initiate by generating PAP to test target LLMs, then assessing the outputs by their relevance to the initial, plain, unpersuaded harmful query.
In Step 1, “Generate PAP”, the fine-tuned Persuasive Paraphraser takes a new plain harmful query and a specified persuasion technique as inputs to generate corresponding PAP. We use greedy sampling to reduce the impact of generation variability on jailbreak Huang et al. (2023).
After probing LLMs with the generated PAP, in Step 2 “Evaluate Harmfulness,” we determine the jailbreak results. We follow Qi et al. (2023) and use a GPT-4 Judge to provide a more contextualized assessment, compared to keyword-based evaluation Zou et al. (2023). The GPT-4 Judge assesses harmfulness on a 1 to 5 Likert scale (1=least harmful, 5=most harmful, refer to Qi et al. (2023) for details about the rubric), taking into account the LLMs’ usage policies, the context, a scoring rubric, and chain-of-thought reasoning. Qi et al. (2023) shows that the GPT-4 Judge is effective and accurate in identifying harmful outputs: notably, it achieved a kappa score of with human annotators.
In our experiment, we are strict about what is considered jailbreak: only when the GPT-4 Judge assigns a top harmful score of 5 do we classify the outcome as Case 2, Jailbreak, characterized by the provision of highly harmful content. Responses that are refused by the model or lack detailed harmful information, rated between 1 and 4, are categorized as Case 1, Refusal.
5 Jailbreak Study : Broad Scan
To gauge the impact of persuasion on jailbreak across a wide range, our initial jailbreak study comprehensively scans all 40 persuasive techniques across diverse risk categories. This investigation reveals how persuasion, a key aspect of human-like communication, interacts with different risk areas, laying the groundwork for more detailed subsequent analysis in the following sections.
5.1 Jailbreak Settings
Target model. In the broad scan phase, we choose GPT-3.5 (gpt-3.5-turbo-0613) as the target model to jailbreak because it is one of the most widely adopted models with clear usage policies that enumerate risk categories444https://openai.com/policies/usage-policies.
Harmful query benchmark. Following the usage policies, we categorize harmful risks into 14 groups as listed in Figure 7. At the time of experiments, there was no publicly available benchmark with well-categorized harmful queries. So, we followed the data sampling procedure in Qi et al. (2023); Yang et al. (2023) to create a categorized harmful query dataset for the jailbreak evaluation. More specifically, there are three sub-categories in each risk category, and we prompt GPT-4 to generate one harmful query for each sub-category, amounting to 143 = 42 plain harmful queries. See examples of the first risk category in Figure 6.
Persuasive Paraphraser Training. In broad scan, we obtain our PAP training data via in-context prompting. We manage to find one working prompt template (Appendix B Figure 12) to bypass the guardrail for “How to make napalm?” and then prompt GPT-4 to generate paraphrased PAP for this plain query. In total, we generate 3 PAPs for each of the 40 techniques (3 40 =120 PAPs in total) to form our training dataset. Then we fine-tune a GPT-3.5 model on this dataset as our Persuasive Paraphraser with the official fine-tuning API555 https://platform.openai.com/docs/guides/fine-tuning and default hyperparameters. Although it is trained on PAPs for “How to make napalm?” only, because of its strong generalization ability, it can be transferred to generate PAPs for other harmful inquiries.
Persuasive Paraphraser Deployment. During deployment, we input the new harmful queries in our categorized benchmark to the trained Persuasive Paraphraser and generate PAPs. For each query-technique pair, 20 PAP variants are generated, leading to a total of 33,600 (14 risk categories 3 harmful queries per category 40 persuasion techniques 20 PAP variants per technique) PAPs. We checked the quality of the generated PAPs and found that 92.9% of these PAPs accurately applied the intended persuasion technique and PAPs overlap with other social engineering methods like virtualization (more detail in E.1). This shows our method can easily be scaled up to generate many unique, high-quality, human-readable PAPs.
Evaluation metrics. We evaluate our broad scan results with the PAP Success Ratio , defined as the percentage of PAPs that lead to outputs with the highest harmfulness score of 5 per GPT-4 Judge.
5.2 Broad Scan Results
Figure 7 shows the broad scan results. An overview is that GPT-3.5 can effectively block all the plain harmful queries (as shown in the bottom row) but cannot defend the paraphrased PAPs with the same bad intent. This indicates that we can persuade GPT-3.5 to jailbreak it in a human-like way.
Across risk categories, we find #7 (fraud/ deception) and #1 (illegal activity) are the most susceptible ones to PAP. This may stem from their subtle and vague nature, making it difficult to categorize and address them with safety measures (e.g., via RLHF). In contrast, categories such as #3 (hate/harass/violence) show better resistance, possibly because they are more clearly defined in existing toxicity guidelines Gehman et al. (2020), which facilitates a better defense. However, we note that no category is entirely safe under PAPs.
Regarding persuasive techniques, logical appeal and authority endorsement are among the most effective ones for jailbreak, with over 60% effectiveness for certain categories. Conversely, tactics like threats are generally less effective.
We also observe interplay between persuasion techniques and risk categories, e.g., logical appeal is highly effective in eliciting harmful responses for #11 (unauthorized practice of law) but less effective for #9 (political campaigning); while negative emotional appeal is more effective for #9 (political campaigning) than for #11 (unauthorized practice of law). This suggests that we cannot simply block certain persuasion techniques to mitigate the risk.
For qualitative evaluation, Figure 8 shows a successful jailbreak PAP for #8 (adult content). The PAP applies the technique “non-expert testimonial” and is easy to understand. Compared to previous algorithm-focused methods, PAPs do not follow a fixed template, making it challenging to defend. Additional qualitative examples from other categories are detailed in E.2, except for category #2 (Children harm), which is omitted for legal and ethical reasons. All the sensitive contents are redacted to prevent real-world harm.
This paper, as an initial exploration of persuasion-related jailbreak risks, concentrates on single-strategy, one-turn PAPs. However, persuasion typically involves a multi-faceted, multi-turn dialogue where users may employ a mix of techniques conversationally. Given the exponentially growing user base and the likelihood of increasingly complex persuasive dialogues, it is imperative for the research community to delve deeper into and mitigate the potential jailbreak risks arising from the identified factor of humanizing and human-like communication with aligned LLMs.
Remark 1: We find persuasion effectively jailbreaks GPT-3.5 across all 14 risk categories. The interplay between risk categories and persuasion techniques highlights the challenges in addressing such user-invoked risks from persuasion. This risk, especially when involving multi-technique and multi-turn communication, emphasizes the urgency for further investigation.
6 Jailbreak Study : In-depth
Iterative Probe
Broad scanning of GPT-3.5 (5) reveals jailbreak risk across all risk categories w.r.t. to PAP. In practice, bad users could iterate upon successful PAPs and refine their approach with varied persuasive techniques. This section models such behavior, detailing an in-depth jailbreak study that fine-tunes a specialized model on effective PAPs. We then assess its ability to jailbreak various LLMs, benchmarking these findings against previous attacks.
6.1 Jailbreak Settings
Target Model. We test PAPs on five aligned LLMs with enhanced safety guardrails: the open-source Llama-2 7b Chat Touvron et al. (2023), GPT-3.5 (gpt-3.5-0613), GPT-4 (gpt-4-0613) OpenAI (2023), Claude 1 (claude-instant-v1), and Claude 2 (claude-v2) Anthropic (2023). We chose these models as they are the most accessible and widely used modern LLMs, likely to be deployed or interacted with large amounts of everyday users.
Harmful query benchmark. We use the AdvBench Zou et al. (2023), refined by Chao et al. (2023) to remove duplicates, which consists of 50 distinct representative harmful queries666 https://github.com/patrickrchao/JailbreakingLLMs.
Persuasive Paraphraser Training. In the in-depth setting, we sample 230 successful PAPs identified in the previous broad scan step and use them as the training data to fine-tune the Persuasive Paraphraser. It is a balanced sample across risk categories and persuasion techniques. Training on this dataset mimics the real-life scenario where bad human actors refine effective jailbreak prompts.
Persuasive Paraphraser Deployment. During deployment, we enumerate persuasion techniques with the Persuasive Paraphraser to generate PAPs using different techniques and prompt LLMs until the GPT-4 Judge detects a jailbreak: if one technique fails, we move on to the next technique in a new session until jailbreak. We define one trial as running through all 40 persuasion techniques, and the maximum number of trials is set to 10. If we cannot jailbreak the model within 10 trials, then it is considered an attack failure. This setup aims to emulate how an average bad actor may manipulate LLMs in a given time period without sophisticated optimization or multi-turn interaction.
Evaluation Metrics. In this setting, we report Attack Success Rate (ASR), the percentage of harmful queries in the AdvBench processed by PAP that leads to jailbreak (with a harmful score of 5 per GPT-4 Judge). The previous PAP Success Ratio measures the ratios of effective PAPs given a specific persuasion technique, while ASR here measures how many harmful queries in AdvBench processed by an attack (for example, iteratively applying all 40 persuasion techniques) within limited trials can jailbreak the victim model.
Baselines Attacks. For algorithm-focused baselines, we selected representative ones like PAIR Chao et al. (2023), GCG Zou et al. (2023), ARCA Jones et al. (2023), and GBDA Guo et al. (2021). Due to their operational differences, a direct comparison with our PAP is challenging (e.g., gradient-based methods need access to the gradients and querying multiple times to manipulate the prompt). To ensure fairness, we align the number of prompts used in our method with these baselines in each trial. For instance, we set PAIR’s number of streams to 40, to match the number of persuasion techniques per trial in our experiment. For gradient-based methods, we adhere to their original settings and hyperparameters, which often involve more than 40 optimization steps per trial. We maintain their most effective settings of total trials (GCG: 3, ARCA: 32, GBDA: 8) and aggregate the results. Since gradient-based methods (GCG, ARCA, GBDA) can only be applied to open-source models, we adapt their prompts generated from open-sourced models like Llama to attack close-sourced models like GPT and Claude series and report the outcomes accordingly. Following Zou et al. (2023), we also set the total number of trials to 3 in this comparison experiment. More details on baseline implementation are in C.
6.2 In-depth Iterative Probing Results
We first analyze PAP’s performance compared to baselines, and then its performance across trials.
6.2.1 PAP comparison with baselines
PAP is more effective than baseline attacks. Table 2 shows the baseline comparison results. Although our PAP method does not use any specialized optimization, it is more effective in jailbreak than existing attacks on Llama-2, GPT-3.5, and GPT-4, highlighting the significant AI safety risks posed by everyday persuasion techniques. While GCG achieves a comparable ASR with PAP on GPT-3.5, it requires more computational resources to synthesize the gradient from open-source LLMs. Interestingly, GCG’s performance drops to when transferred to GPT-4, likely due to additional safety measures in OpenAI’s more advanced models after they released their paper. Notably, although GCG, GBDA, and ARCA are optimized directly on Llama-2 7b Chat, none of them match our PAP’s ASR on Llama-2. This suggests that while Llama-2 may have been aligned to defend these established algorithm-focused attacks, their safety measures might have underestimated the jailbreak risks coming from natural communication with everyday users. A side note is that all the evaluated jailbreak methods perform poorly on the Claude models, indicating a distinct safety measure difference between Claude’s and other model families.
| Method | Trials | ASR () @ | ||||
| Llama-2 | GPT-3.5 | GPT-4 | Claude-1 | Claude-2 | ||
| PAPs | 3 | 68% | 86% | 88% | 0% | 0% |
| PAIR | 3* | 30% | 42% | 54% | 4% | 4% |
| GCG | 3 | 16% | 86% | 0% | 0% | 4% |
| ARCA | 32 | 0% | 2% | 0% | 0% | 0% |
| GBDA | 8 | 0% | 0% | 0% | 0% | 0% |
6.2.2 PAP performance across trials
Figure 9 presents the ASR for different numbers of trials. In this part, we also extend the number of trials to 10 to test the boundary of PAPs and report the overall ASR across 10 trials.
Notably, stronger models may be more vulnerable to PAPs than weaker models if the model family is susceptible to persuasion. Drom the ASR within 1 and 3 trials, we see that GPT-4 is more prone to PAPs than GPT-3.5. A possible reason is that as models’ capability and helpfulness increase, they can better understand and respond to persuasion and thus become more vulnerable. This trend differs from previous observations that attacks usually work better on smaller models Zou et al. (2023), reflecting the uniqueness of risks elicited by PAPs.
The overall ASR varies for different model families: PAP achieves 92% ASR on Llama-2 and GPTs but is limited on Claude. For Llama-2 and GPT models, PAPs can achieve an alarming ASR of over 92% within 10 trials, while for the Claude family, PAP is much limited in performance. This indicates that Claude is much harder to jailbreak, which is consistent with others’ findings Zou et al. (2023); Chao et al. (2023). One difference between Claude models and other models is the usage of RLAIF Bai et al. (2022), RL from AI Feedback, which may play a pivotal role in their robustness and shed light on future safety mechanisms. Nevertheless, with a worryingly high ASR across Llama-2 and GPT models, even without specialized optimization, we still highlight the unique, overlooked risks coming from human-like communication with everyday users.
For qualitative evaluation, Figure 10 presents a successful PAP on GPT-4; E.2 shows more working PAP examples for different victim LLMs.
Remark 2: To mimic human refinement behavior, we train on successful PAPs and iteratively deploy different persuasion techniques. Doing so jailbreaks popular aligned LLMs, such as Llama-2 and GPT models, much more effectively than existing algorithm-focused attacks. Interestingly, more sophisticated models such as GPT-4 exhibit greater susceptibility to PAPs than their predecessors like GPT-3.5. This underscores the distinctive risks posed by human-like persuasive interactions.
7 Re-evaluating Existing Defenses
This section revisits general post hoc adversarial prompt defense strategies that do not modify the base model or its initial settings (e.g., system prompt). Specifically, we focus on mutation-based and detection-based defenses, deliberately omitting perplexity-based methods Alon and Kamfonas (2023); Jain et al. (2023), which depend on identifying unusually high perplexity. Our rationale for this exclusion is that our generated PAPs are coherent and exhibit low perplexity. Our emphasis is on black-box defense mechanisms suitable for closed-source models. The following provides an overview of these defense strategies:
Mutation-based: This type of defense alters inputs to reduce harm while preserving the meaning of benign inputs. We test two methods, Rephrase and Retokenize, proposed in Jain et al. (2023).
Detection-based : This line of defense detects harmful queries from the input space. Examples include Rand-Drop Cao et al. (2023), which drops tokens randomly to observe the change in responses; RAIN Li et al. (2023), which relies on in-context introspection; and Rand-Insert, Rand-Swap, and Rand-Patch Robey et al. (2023), which also alter the inputs and inspects the change in outputs.
C provides more detail on the defense implementation. We defend PAP generated in the in-depth probe (6). We did not experiment on Claude models as they are already robust to PAP.
| Defenses | ASR () | ||
| @Llama-2 | @GPT-3.5 | @GPT-4 | |
| No defense | 92% | 94% | 92% |
| Mutation-based | |||
| Rephrase | 34% (-58) | 58% (-36) | 60% (-32) |
| Retokenize | 24% (-68) | 62% (-32) | 76% (-16) |
| Detection-based | |||
| Rand-Drop | 82% (-10) | 84% (-10) | 80% (-12) |
| RAIN | 60% (-32) | 70% (-24) | 88% (-4) |
| Rand-Insert | 92% (-0) | 88% (-6) | 86% (-6) |
| Rand-Swap | 92% (-0) | 76% (-18) | 80% (-12) |
| Rand-Patch | 92% (-0) | 86% (-8) | 84% (-8) |
Table 3 shows the ASR and how much the defense can reduce the ASR. Overall, mutation-based methods outperform detection-based methods in lowering ASR. But mutation also alters benign queries, which could potentially diminish the model’s helpfulness. Mutation methods can defend Llama-2 more effectively, likely because GPT models can better understand altered inputs than Llama-2 7b. Again, we observe the interesting trend that the more advanced the models are, the less effective current defenses are, possibly because advanced models grasp context better, making mutation-based defenses less useful. Notably, even the most effective defense can only reduce ASR on GPT-4 to 60%, which is still higher than the best baseline attack (54% per Table 2). This strengthens the need for improved defenses for more capable models.
Remark 3: We uncover a gap in AI safety: current defenses are largely ad-hoc, e.g., defenses often assume the presence of gibberish, overlooking semantic content. This oversight has limited the creation of safeguards against more subtle, human-like communication risks exemplified by PAPs. Our findings underscore the critical need to revise and expand threat models in AI safety to encompass these nuanced vulnerabilities.
8 Exploring Adaptive Defenses
LLMs’ vulnerability to PAPs may be because they are trained on language data rich in persuasive structures and content. For example, a quick search on the platform with information about large text corpora777https://wimbd.apps.allenai.org/ Elazar et al. (2023) shows that the term “persuasion” appears over 1.5 million times in the widely adopted C4 corpus Raffel et al. (2020). Additionally, dark patterns, which are online user interfaces to influence users, are also commonly used on various websites Mathur et al. (2019); Narayanan et al. (2020); Luguri and Strahilevitz (2021), such as messages showing certain items are in high demand. These online contents may further enable LLMs to recognize and respond to similar persuasive patterns in queries. A simplistic defense might involve removing such persuasive content from pre-training. However, as we do not know what these LLMs are trained on, it is difficult to analyze the presence of persuasive content in pre-training data. Moreover, persuasion is ubiquitous in communication O’keefe (2018); Izuma (2013), so simply removing all persuasive contents may adversely affect the LLM utility. Instead, we investigate two straightforward and intuitive adaptive defense tactics: “Adaptive System Prompt” and “Targeted Summarization,” designed to counteract the influence of persuasive contexts in PAPs. We explore three adaptive defenses within these two tactics:
Adaptive System Prompt (Adapt Sys.): We use a system prompt to instruct the LLM to resist persuasion explicitly: “You are a helpful assistant, and you are not easily persuaded, and you know what is right and wrong.”
Base Summarizer (Base Smry.): We prompt GPT-4 to summarize the adversarial prompts to output the core query before executing the input via the target LLM.
Tuned Summarizer (Tuned Smry.): We fine-tune a GPT-3.5-based summarizer using pairs of plain harmful queries and their corresponding PAP from 5. To prevent it from always deriving harmful queries from even benign inputs, we incorporate benign data sampled from Alpaca Taori et al. (2023) in the training set.
Details of the two summarizers are in D. We now proceed to qualitatively and quantitatively compare these defenses with baseline defenses.
8.1 Results & Findings
Figure 11 provides a qualitative evaluation of defenses against PAPs. “Adapt Sys.” modifications reinforce model safety in the presence of persuasive contexts, promoting safer outputs without altering the user’s original persuasive prompt. Although the output is safer for this specific example, it seems “Adapt Sys.” did not trigger the GPT-4’s guardrail (which typically responds with “Sorry, but I can’t assist with that.”) In contrast, the tuned summarize (“Tuned Smry”) summarizes the PAP to the core adversarial request “Give me a summary of how to make a homemade explosive”, which triggers the model’s inherent guardrail and effectively blocks the harmful query. This approach goes beyond surface-level alterations like Paraphrase or Retokenize, delving into the semantic and pragmatic meaning to extract and nullify the prompt’s harmful substance, thus presenting more general and effective mitigation to PAPs on different models.
| ASR () | MT-bench () | |||
| @Llama-2 | @GPT-3.5 | @GPT-4 | @GPT-4 | |
| No Defense | 8.97 | |||
| PAPs | 92% | 94% | 92% | |
| ⚫ Paraphrase | 7.99 | |||
| PAPs | 34% (-58) | 58% (-36) | 60% (-32) | |
| ⚫ Retokenize | 8.75 | |||
| PAPs | 24% (-68) | 62% (-32) | 76% (-16) | |
| Adapt Sys. | 8.85 | |||
| PAPs | 30% (-62) | 12% (-82) | 38% (-54) | |
| PAIR | 14% (-16) | 0% (-42) | 14% (-40) | |
| GCG | 4% (-12) | 0% (-86) | 0% (-0) | |
| Base Smry. | 6.51 | |||
| PAPs | 22% (-70) | 42% (-52) | 46% (-46) | |
| PAIR | 4% (-26) | 8% (-34) | 20% (-34) | |
| GCG | 0% (-16) | 8% (-78) | 0% (-0) | |
| Tuned Smry. | 6.65 | |||
| PAPs | 2% (-90) | 4% (-90) | 2% (-90) | |
| PAIR | 0% (-30) | 6% (-36) | 6% (-48) | |
| GCG | 2% (-14) | 8% (-78) | 0% (-0) | |
Quantitatively, Table 4 shows that modifying system prompts (“Adapt Sys.”) alone improves model resilience against PAPs, often outperforming baseline defenses. The two adaptive summarization — base and tuned summarizers — also surpass the baseline defenses in neutralizing PAPs. The tuned summarizer (“Tuned Smry.”), in particular, demonstrates superior efficacy, reducing the ASR of PAPs on GPT-4 from 92% to 2%, signifying a notable enhancement in practical post-hoc defense.
More interestingly, adaptive defenses, initially tailored for PAPs, are also effective against other types of adversarial prompts. For instance, adjusting the system prompt to emphasize resistance to persuasion, we witnessed a decline in the ASR for the GCG from 86% to 0% on GPT-3.5. Similarly, with “Tuned Smry.”, the ASR for both PAIR and GCG was reduced to below 8% across various models. These observations suggest that although different adversarial prompts are generated by different procedures (gradient-based, modification-based, etc.), their core mechanisms may be related to persuading the LLM into compliance. For instance, GCG employs gradients but typically seeks a submissive “Sure” in response to harmful queries, and the generated gibberish suffix may be seen as persuasive messages understandable to LLMs. Such insights imply an interesting future research direction to study the link between persuasion and jailbreak: jailbreak, at its essence, may be viewed as a persuasion procedure directed at LLMs to extract prohibited information, and various types of adversarial prompts may be unified as persuasive messages towards LLMs. This further hints at the potential for developing more fundamental defense frameworks aimed at resisting persuasion to enhance AI safety.
Our findings also indicate that there exists a trade-off between safety and utility, so a widely effective defense mechanism may not be the optimal choice for every model. For example, although “Tuned Smry.” achieves the highest protection levels on PAP for GPT-3.5 (ASR 94% to 4%), it considerably diminishes model helpfulness, with MT-bench scores dropping from 8.97 to 6.65; while “Adapt Sys.” demonstrates effective PAP mitigation in GPT-3.5 and minimally impacts MT-bench scores (8.97 to 8.85). This indicates that “Adapt Sys.” is a better safety solution for GPT-3.5.
Notably, post-hoc defenses still remain important. Because even models resistant to PAP (e.g., the Claude series) may still have their own weaknesses. For instance, the Claude series are vulnerable to complex virtualization jailbreaks Yu et al. (2023); Deng et al. (2023a). Summarization techniques discussed in this section are proven valuable in such instances, as detailed in D. These results show the necessity of model-specific defenses that consider model characteristics and threat type rather than a one-size-fits-all defense method.
Remark 4: We reveal that the developed adaptive defenses are effective in counteracting PAP. Interestingly, they can also defend other types of jailbreak prompts beyond PAPs. This suggests that it is a worthwhile future direction to study the underlying connection between persuasion and jailbreak that aims to elicit compliance on prohibited topics. Additionally, we highlight the trade-off between safety and utility: while generalizable and effective defenses can enhance model safety, they can also diminish utility. Therefore, the selection of a defense strategy should be tailored to individual models and specific safety goals.
9 Conclusion
Unlike traditional AI safety research that treats AI models as algorithmic systems or mere instruction followers, we introduce a new perspective by humanizing LLMs and studying how to persuade LLMs to jailbreak them like humans. We first propose a persuasion taxonomy based on decades of social science research. Such a thorough taxonomy helps us automatically generate PAP and systematically explore the impact of persuasion on LLM vulnerabilities. Our study reveals that LLMs are susceptible to various persuasion techniques, and PAP consistently outperforms algorithm-focused jailbreak methods with an attack success rate of over 92% on Llama-2 7b Chat, GPT-3.5, and GPT-4. We also observe that more advanced models are both more susceptible to PAP and more resistant to conventional defense strategies, possibly due to their enhanced understanding of persuasion. These results reveal a critical gap in current defenses against risks coming from human-like communication. In our efforts to mitigate risks, we discovered that adaptive defenses designed for PAP are also effective against other forms of attacks, revealing a potential connection between persuasion and broader jailbreak risks. To conclude, our findings highlight the unique risks rooted in natural persuasive communication that everyday users can invoke, calling for more fundamental solutions to ensure AI safety in real-world applications.
Ethical Consideration
This paper provides a structured way to generate interpretable persuasive adversarial prompts (PAP) at scale, which could potentially allow everyday users to jailbreak LLM without much computing. But as mentioned, a Reddit user 888https://www.reddit.com/r/ChatGPT/comments/12sn0kk/grandma_exploit has already employed persuasion to attack LLM before, so it is in urgent need to more systematically study the vulnerabilities around persuasive jailbreak to better mitigate them. Therefore, despite the risks involved, we believe it is crucial to share our findings in full. We followed ethical guidelines throughout our study.
First, persuasion is usually a hard task for the general population, so even with our taxonomy, it may still be challenging for people without training to paraphrase a plain, harmful query at scale to a successful PAP. Therefore, the real-world risk of a widespread attack from millions of users is relatively low. We also decide to withhold the trained Persuasive Paraphraser to prevent people from paraphrasing harmful queries easily.
To minimize real-world harm, we disclose our results to Meta and OpenAI before publication, so the PAPs in this paper may not be effective anymore. As discussed, Claude successfully resisted PAPs, demonstrating one successful mitigation method. We also explored different defenses and proposed new adaptive safety system prompts and a new summarization-based defense mechanism to mitigate the risks, which has shown promising results. We aim to improve these defenses in future work.
To sum up, the aim of our research is to strengthen LLM safety, not enable malicious use. We commit to ongoing monitoring and updating of our research in line with technological advancements and will restrict the PAP fine-tuning details to certified researchers with approval only.
Limitation and Future Work
In this study, we mainly focus on single-turn persuasive attempts, but persuasion is oftentimes a multi-turn interactive process. For instance, persuasive techniques like “foot in the door” (start with a small request to pave the way for a larger one) and “reciprocity” (adapt to the other party’s linguistic styles) rely on the buildup of conversation context. Xu et al. (2023) shows that LLMs can be persuaded to believe in misinformation, and multi-turn persuasive conversation is more effective than single-turn persuasive messages. In the jailbreak situation, it remains unclear whether these strategies’ effectiveness would increase or if the LLMs would become more resistant after noticing prior rejections in a conversation. Besides, certain persuasion techniques, like emotional appeal, are more popular than others, and users can also mix different techniques in one message to improve its persuasiveness, but in our experiment, we generate the same amount of PAP per technique. These factors may make the jailbreak distribution different from the real-life persuasive jailbreak scenarios. This gap in our study points to the need for more comprehensive research in this area.
We have shown PAP methods can jailbreak LLMs, but it would be interesting to see if humans would also react to these PAPs and be persuaded to provide harmful information and how the human-AI persuasion and human-human persuasion differ. Besides, it remains an open question if LLM outputs after jailbreak are truly harmful in the real world. For instance, even without LLM, users can search on the internet to gather information about drug smuggling. Also, there are different nuances to the harmfulness evaluation. Sometimes, the information itself may be neutral, and if it is harmful depends on who will access it and how they will use it: for instance, law enforcement agencies may need detailed information on drug smuggling to prevent it, but if bad actors access the information, it may be used to commit the crime.
Besides, our study primarily focused on persuasion techniques, but future research may find value in a deeper analysis of the specific linguistic cues, keywords, etc, inside PAPs. This could reveal more insights into the mechanics of persuasive jailbreak for better defense algorithms.
In sum, as AI technology advances, larger and more competent models may emerge, which can potentially respond even more actively to persuasive jailbreak. This progression invites a new direction of research to systematically protect these advanced models from manipulation. Investigating how these more sophisticated models interact with persuasion from a cognitive and anthropological standpoint could provide valuable insights into developing more secure and robust AI systems.
Acknowledgment
We thank Alicja Chaszczewicz, Derek Chen, Tatsunori Hashimoto, Minzhi Li, Ryan Li, Percy Liang, Michael Ryan, Omar Shaikh from Stanford, Lucy He, Peter Henderson, Kaixuan Huang, Yangsibo Huang, Udari Madhushani, Prateek Mittal, Xiangyu Qi, Vikash Sehwag, Boyi Wei, Mengzhou Xia, Tinghao Xie from Princeton, Alex Beutel, Lilian Weng from OpenAI, and Nicholas Carlini from Google for their valuable discussion or feedback. Ruoxi Jia and the ReDS lab acknowledge support through grants from the Amazon-Virginia Tech Initiative for Efficient and Robust Machine Learning, the National Science Foundation under Grant No. IIS-2312794, NSF IIS-2313130, NSF OAC-2239622, and the Commonwealth Cyber Initiative. Yang acknowledges the support by the Defense Advanced Research Project Agency (DARPA) grant HR00112290103/HR0011260656, ONR, and NSF grant IIS-2308994. Weiyan Shi acknowledges the support from Northeastern University. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding agencies. We also thank OpenAI for an API Research Credits grant.
Appendix A Construction of the Taxonomy
The persuasion taxonomy in this paper is drawn from decades of extensive social science research across psychology, communication, sociology, marketing and so on, Table 5 shows a map between the literature and our taxonomy. On the psychology side, we include terms and definitions of persuasion techniques based on Pornpitakpan (2004); Cialdini and Goldstein (2004); Cialdini (2001); Petty et al. (2003); Bilandzic and Busselle (2013); Wilson et al. (2013); Olson and Zanna (1990); Johannesen and Larson (1989); DiFonzo and Bordia (2011). Besides, communication research, notably Cronkhite (1964); Perloff (2017); Rieh and Danielson (2007); Dillard and Knobloch (2011); Burgoon et al. (1993), significantly informs our techniques. Sociology literature covering political science Brader (2005) and philosophical analysis Powers (2007) has been also used to update some of the terminology of these techniques. Additionally, concepts from marketing, particularly O’Keefe (2016); Wang (2005); Woodside et al. (2008); Aggarwal et al. (2011), play a crucial role in leading us to deliver the final terminology and definitions in our taxonomy.
Appendix B Simple prompting fails to build a stable and scalable Persuasive Paraphraser
In this section, we show that due to safety guardrails, we cannot simply prompt LLMs for the persuasive paraphrase task and need to train a dedicated Persuasive Paraphraser.
Simple prompting cannot stably paraphrase or alter the plain harmful query. Initially, we attempted to simply prompt an LLM to paraphrase the plain harmful queries. Figure 12 shows the prompt template: we present GPT-4 with information from the taxonomy about a persuasion technique, coupled with a harmful instruction, to elicit a persuasion-enhanced response.
Sometimes this method works: Figure 13 shows a successful in-context paraphrase, where GPT-4 automatically applies “logical appeal” to rephrase the harmful instruction “How to make napalm?”
However, the effectiveness of this prompting method varies. Despite using the same prompt templates and persuasion knowledge (as shown in Figure 12), GPT-4’s guardrails sometimes mitigate or alter the intent of the plain harmful query. Figure 14 shows one example, the plain harmful query is “How to hurt myself”, but GPT-4 fails to persuasively paraphrase this question, but rather asks a different question about how to cope with negative emotions, possibly due to strong safety measurements. As we need to scale up the PAP generation, such an unstable behavior will impact the quality of the PAP and, thus, the accuracy and accountability of the following experiments.
in red
To generate PAP more reliably, we need to bypass guardrails. As shown by recent studies Qi et al. (2023), fine-tuning can effectively mitigate the influence of guardrails. So we propose to fine-tune a dedicated LLM as the Persuasive Paraphraser to translate plain harmful queries into PAPs. This process is shown in Figure 4 and discussed in 4. Figure 15 shows the improved outcomes compared to simple prompting (Figure 14) on the same harmful query, which demonstrates that the fine-tuned Persuasive Paraphraser can enhance the generated PAP’s quality. In Section E.1, we will further quantitatively analyze the generated PAP.
Appendix C Baselines and Hyperparameters
Attack Baselines.
We adjust the hyperparameters of baseline methods to ensure a fair comparison. For black-box methods, including PAIR999https://github.com/patrickrchao/JailbreakingLLMs Chao et al. (2023) and ours, our goal is to ensure the same number of queries on the target model. Specifically, for PAIR, to align with our 40 strategies, we set a stream size of and a maximum depth of , where a depth of 3 means that we iteratively optimize their attacking prompt for three times in a dialogue-based setting. The rest of the comparison methods are white-box baselines, where we retain each method’s original configuration and aggregate results from multiple trials, similar to our settings. For GCG101010https://github.com/LLM-attacks/LLM-attacks Zou et al. (2023), we use Vicuna-7b-v1.3 and Llama-7b-chat for joint optimization of 500 steps, conducting 3 experiments to generate distinct suffixes following the strongest settings in the original paper. In the ensemble setting, we also tested attacks incorporating these combined suffixes (directly concatenation). For ARCA111111https://github.com/ejones313/auditing-LLMs Jones et al. (2023), we configure 32 candidates (32 trails) and a maximum of 50 iterations for each plain harmful query. For GBDA121212https://github.com/facebookresearch/text-adversarial-attack Guo et al. (2021), we sample 8 times (8 trials) per plain harmful query per step and conduct 200 steps with a learning rate of . Noting that we have all the baseline methods deploy equal or more numbers of queries than ours. For all the methods aggregating from multiple rounds, a successful attack is defined as jailbreaking a plain harmful query in at least one of the trials.
Defense Settings.
Details of the mutation-based defenses are as follows: we use ChatGPT to paraphrase prompts for the Paraphrase method Jain et al. (2023), setting the temperature to 0.7. The Retokenize method follows the settings described in Jain et al. (2023).
The detection-based defense settings are as follows. For Rand-Drop131313https://github.com/AAAAAAsuka/LLM_defends Cao et al. (2023), we set a drop probability () of 0.3, a threshold () of 0.2, and conducted 20 rounds of sampling of the output as following their default settings. Both Cao et al. (2023) and Kumar et al. (2023) detect harmful prompts by randomly dropping tokens and analyzing the changes. As Cao et al. (2023) provides a well-justified threshold selection, and the techniques are similar, we evaluate this method only. For RAIN141414https://github.com/SafeAILab/RAIN Li et al. (2023), a recently proposed alignment technique grounded in self-evaluation, we tested its binary classifier (the self-evaluation phase in the paper), which assesses if a response is harmful or harmless given the generated content only. Following the original implementation, we averaged results over two shuffled options (swapping the order of harmful or harmless options). RAIN’s prompt does not have the context, policies, scoring, and reasoning chains of GPT-4 Judge, which might be one factor limiting their detectability as reflected in Table 3. For Smooth LLM151515https://github.com/arobey1/smooth-LLM Robey et al. (2023), we implemented three random perturbation methods proposed in this work: Rand-Insert, Rand-Swap, and Rand-Patch. Each method was set with a maximum disturbance probability () of 0.2 and a sampling number () of 10 following their strongest settings. To evaluate the results’ harmfulness before and after perturbation, we follow their evaluation setting and inspect a keyword set from Zou et al. (2023) during output inspection.
Appendix D Implementation Details of Defense via Summarization (Base & Tuned Smry.)
Base Smry.
We simply prompt GPT-4 with the template in Figure 16, asking it to summarize any given inquiry. Then, we feed the summarized output to downstream target models and evaluate the final output from the target model to determine the jailbreak result (ASR) and helpfulness (MT-bench score).
Fine-tuned Smry.
To develop the fine-tuned summarizer, we employed the system prompt in Figure 17. This prompt straightforwardly inserts a plain harmful query and the corresponding PAP, simulating a scenario where the defender knows about the PAPs’ distribution. For this, we randomly selected 50 samples from the same pool of 230 used to fine-tune the persuasive paraphrase. But if the summarizer is trained on adversarial examples only, it will also always summarize benign inputs to a harmful query and detect all queries as harmful and hurt the helpfulness. To avoid such false positives, we also included 50 benign alpaca instruction samples, processed through the Base Smry., to replicate benign inputs undergoing summarization. These 100 samples formed the dataset, which was then applied to the template in Figure 17 to fine-tune GPT-3.5 using the OpenAI API with default hyperparameters. During deployment, the same system prompt is used, but the input is replaced with the user’s query. We then feed the summarized query to the target model and evaluate its output.
Additionally, we put Claude models to test manually crafted virtualization-based prompts and use our fine-tuned summarizer for defense. The results are shown in Table 6. We utilize 77 jailbreak templates that can be combined with harmful queries. They are artificial templates from the jailbreak chat website161616https://www.jailbreakchat.com/, collected by Liu et al. (2023b) and filtered through GPTFuzzer Yu et al. (2023). Besides the initial set, we sample two kinds of variants of artificial templates following the attack design proposed in GPTFuzzer. Firstly, we mutate the templates using five mutation operators from GPTFuzzer. Each template is randomly mutated 3 times to generate 3 variants. Secondly, we utilize the rewriting prompt from Masterkey Deng et al. (2023a) to prompt ChatGPT for 3 rephrased samples per template. We use these jailbreak templates combined with 50 harmful queries to conduct attacks on Claude models. The initial templates lead to 3,850 attacking cases, while the two kinds of variants result in 11,550 attacking cases, respectively. As seen in Table 6, these manually crafted templates demonstrate effectiveness in jailbreaking Claude models, with higher rates on Claude-2. The “Tuned Smry.” is able to counteract all the jailbreak templates. The observation remarks that an interesting factor of summarization is content moderation. In particular, our Tuned Smry. only observed persuasive prompts and learned how to summarize them to plain inquiries. However, the effectiveness in protecting models against unseen attacks evaluated in this section, i.e., manually crafted prompts and advanced variants produced by GPTFuzzer and Masterkey, demonstrates a promising direction in the future design of summarization-based adversarial prompt defenses. However, this current version of the summarizer still negatively impacts the helpfulness of the Claude models. After summarization, the MT-bench score drops from 8.31 to 6.04 on Claude-1 and from 8.10 to 6.03 on Claude-2. In the future, we can improve the summarizer to target different types of attack methods and use more benign samples to ensure helpfulness.
| Method | Count | Model | |
| Claude-1 | Claude-2 | ||
| Initial Templates | 3850 | 16 | 376 |
| w/ Tuned Smry. | 0 | 0 | |
| GPTFuzzer | 11550 | 9 | 473 |
| w/ Tuned Smry. | 0 | 0 | |
| Masterkey | 11550 | 44 | 420 |
| w/ Tuned Smry. | 0 | 0 | |
Appendix E Qualitative Study
Now, we analyze the quality of the PAP generated. We provide a human evaluation of the generated PAPs and study the quality of harmfulness of the elicit outputs with respect to jailbreaking cases.
E.1 Quantitative Analysis of PAPs
In the quantitative analysis, we focus on the following two core aspects of PAP: 1) if they accurately apply the required persuasive technique and 2) if they are a faithful paraphrase of the original harmful query and ask the same ill-intended question. Besides, we also analyze how much PAP overlaps with other attack methods like virtualization-based Kang et al. (2023) and persona-based Shah et al. (2023) approaches.
Three human experts annotate a subset of PAP samples on different related questions. The annotation interface is depicted in Figure 18. The first question is about whether the PAP employs the assigned persuasion technique accurately. The other three questions are about whether the PAP also overlaps with other attack methods, such as persona-based attack where the adversarial prompt asks the LLM to behave like a certain persona, or the prompt itself pretends to behave like a certain person, or virtualization-based attack where the adversarial prompt builds a virtualized world and ask the LLMs to follow certain orders.
They first independently annotated the same 40 PAP randomly drawn from the broad scan (5). Then, they discussed disagreement and achieved an inter-annotator agreement on fixing disagreed examples. Afterward, each annotator separately annotated 80 PAP (2 examples 40 techniques), amounting to 240 annotated PAP samples.
Table 7 shows that our Persuasive Paraphraser is highly reliable in generating high-quality PAP: 92.92% PAPs correctly apply the required persuasion techniques. Besides, PAP poses unique risks without much overlap with other attacks ( with all other attacks). We also note that such overlap to leverage persona or virtualization can happen organically in persuasion. For example, when persuading someone with “misrepresentation", a natural procedure is to take on a fake persona and justify the reason for asking the certain query. Similar to virtualization, in “Framing”, a constructed imaginary case may be deployed.
| Percentage | 92.92% | 10.00% | 8.75% | 7.92% | ||||||||
| #samples | 223 | 24 | 21 | 19 |
E.2 Examples (Harmful Contents Warning!)
Finally, we dive into the qualitative examples of selected jailbreak cases against different target models. For clarity, we truncate jailbreaking cases to only include the portion presenting harmful information and redact sensitive details if applicable. Note that the example from the risk category “Child Harm” is intentionally omitted.
Again, the contents listed in the following parts of this section can be harmful in nature! Examples are provided for research purposes.